# Gaussian Splatting in Style

### v1

@@ Abstract
Scene stylization extends the work of neural style transfer to three spatial dimensions. A vital challenge in this problem is to maintain the uniformity of the stylized appearance across a multi-view setting. A vast majority of the previous works achieve this by optimizing the scene with a specific style image. In contrast, we propose a novel architecture trained on a collection of style images, that at test time produces high quality stylized novel views. Our work builds up on the framework of 3D Gaussian splatting. For a given scene, we take the pretrained Gaussians and process them using a multi resolution hash grid and a tiny MLP to obtain the conditional stylised views. The explicit nature of 3D Gaussians give us inherent advantages over NeRF-based methods including geometric consistency, along with having a fast training and rendering regime. This enables our method to be useful for vast practical use cases such as in augmented or virtual reality applications. Through our experiments, we show our methods achieve state-of-the-art performance with superior visual quality on various indoor and outdoor real-world data.
@@
Figure 1: Given multi-view images of a real-world scene, we perform the task of scene stylization. We train a network that, provided with style image at test time, generates stylized novel views of the scene conditioned on the input style in real-time that are consistent in 3D space. In contrary to popular scene stylization approaches of fitting a scene with each new style, we learn this mapping via a neural network and therefore, can generate novel views of the scene even for unseen styles at a rate of roughly 150FPS.
Figure 2: The motivation from our work stems from the requirement of a specialized method that while stylizing a scene considers the spatial information into account. We show that to generate stylized novel views of a scene, it is not sufficient to stylize the rendered views or train a scene representation model on stylized 2D images. It leads to loss of information such as deformity in solid truck’s body shown above.
@@ 1 Introduction
Driven by the curiosity of how a particular image would appear if painted by Van Gogh, Picasso or any artist, style transfer has been at the heart of the image processing community for a long time. It can be seen as a challenge of texture transfer. Earlier works of [13, 12, 81, 37] devised powerful algorithms for texture synthesis. Further advancements in deep learning gave rise to works such as [16, 31, 27] that approach the problem by suitably trained neural networks. The core idea of neural style transfer is to apply the features of a style image such as colors, texture, and aesthetics to the content image while keeping information inside the content image intact. The ability of neural networks to grasp the features of an image at multiple hierarchical levels made them a preferred choice over non-learning based algorithms[16]. Fuelled by the availability of 3D data, this phenomenon of style transfer was extended to the stylization of scenes[28, 26, 43, 87], and is referred to as scene stylization.
@@
In this work, we propose a solution for the task of scene stylization given a style image. The ability to control the visual properties of a scene is central to many practical applications such as avatar modeling or scene editing [60, 51, 64, 91] to name a few. With the wide adaptability of various augmented and virtual reality appliances gaining momentum, it is not only necessary that these techniques are correct, but also need to execute in real-time for maximum consumer satisfaction.
@@
Over the past few years, various different methods have been proposed to efficiently represent a complex 3D scene. [59, 82, 77, 71] were some of the methods leveraging an explicit pointcloud representation. The domain witnessed an incredible pace of growth following the pioneering work of Neural Radiance Fields(NeRFs)[50, 49]. NeRFs gained large popularity, primarily due to its ability to store the information of a complex scene as the weights of a multi-layer perceptron(MLP), while maintaining great detailing in the resulting novel view synthesis. One drawback of radiance fields was the slow speed of training and at test time. This was followed up by many variants of NeRFs such as [66, 86, 1, 54, 57, 3] that improved the seminal work significantly. Recently, the work of 3D Gaussian Splatting(3DGS)[33] further pushed the boundaries of novel view synthesis, especially in terms of consistency and speed. 3DGS is, by design, more explicit in terms of storing scene information in the form of 3D Gaussians, with each Gaussian having certain properties attached to it. Additionally, 3DGS renders views in real-time, thus making it highly suitable for practical applications.
@@
In this paper, we tackle the problem of scene stylization as that of stylizing a set of pre-trained Gaussians given a set of style images. Therefore, we name our approach as Gaussian Splatting in Style(GSS). We then use these Gaussians conditioned on a style image to obtain stylized views of the complex 3D scene. Utilizing these Gaussians as the backbone of our approach ensures spatial consistency. Furthermore, our approach does not add any overhead to the existing real-time rendering speed of 3DGS, making it a preferred choice for applications in AR/VR ecosystems. We support our method by obtaining the most consistent short-term and long-term views as a quantitative metric. To further strengthen our approach, we also compare the superior quality of stylized novel views generated by our method in a qualitative fashion against other methods. To summarise, our contribution in this paper are as follows:
@@
  • We develop a novel state-of-the-art method, GSS, to perform neural scene stylization in real-time based on 3D Gaussian splatting. To the best of our knowledge, we are the first method to perform scene stylization using 3D Gaussians.
@@ 2.2 Novel View Synthesis
2.3 Style Transfer
@@
Neural style transfer is the task of transferring the style of a source image onto the target(or content) image[16]. It is a complex task as the transfer needs to preserve the information present in the target image. This problem can also be formulated as that of texture transfer, and was tackled in the works of [13, 12, 81, 37] before the widespread adoption of neural networks. Later, the pioneering work of [16] successfully applied the use of neural networks to the task of style transfer. The work made use of a pretrained VGG[72] to extract the semantic information from an image at different hierarchical levels. [16] optimized a loss function composed of two terms, content loss and style loss, which represented the similarity of the generated image with the content image and the style image respectively. The style loss was optimized by using the Gram matrix. The primary objective of Gram matrix was to encapsulate the correlations between the different features. Since [16] worked by running an optimization for each target and style image, it was not suitable for real-time applications. This was further improved by [40, 31, 78, 27] by being real-time and providing the flexibility of general style conditioning. Additionally, the works of [44, 30] added the sophistication of monocular depth estimation for preserving the depth information present in the source image. Furthermore, works of [25, 19, 15, 4, 6, 80] extended the work of 2D style transfer to a video sequence by making use of advancements in methods like optical flow[11]. However, [26] and our experiments show that simply extending 2D style transfer to 3D scenes result in visual artifacts such as blurriness and inconsistency across different views, as also depicted in Figure 2.
2.4 Scene Stylization
@@
Stylizing scenes given a style image has gained prominence in recent years, especially with the arrival of AR/VR headsets in the mainstream. One way to distinguish between different methods is based on their way of scene representation. Works of [26, 52, 2] work with a point cloud based representation. [26] for instance, project the style features from 2D into 3D and transform them into that of the style image before re-projecting them back into 2D for getting stylized views. Similarly, [23] performs scene stylization of indoor scenes in the form of a mesh. Recently, radiance fields [50] gained huge popularity due to their ability to generate novel views with a very high quality. Due to this, a lot of works made use of a NeRF backbone to perform scene stylization. One direction of approaching the problem is by focusing to optimize a radiance for each given style image. Methods such as [32, 87, 41, 8, 56, 84, 89, 61, 90] follow this outline. Methods such as [7, 8] make use of a hypernetwork to impart the style image features onto the rendered views of a complex 3D scene. On the other hand, works such as [28, 43] provide a generalizable scene stylization framework, i.e a stylized scene can be generated at inference time given a conditioned style image input. Our work also follows in this category. Another direction in this domain lies making use of additional information such as depth maps [32] since their primary focus lies in geometry preservation. In this work, instead of a radiance field, we make use of a Gaussian splatting[33] framework. To the best of our knowledge, we are the first method that builds up on a 3DGS backbone for image-based scene stylization.
@@ 3 Preliminaries
3D Gaussian Splatting(3DGS)[33] is an explicit method for scene representation, that is especially useful for high quality real-time rendering. It can be considered as a form based on point cloud representation. Each scene in 3DGS is represented as a bunch of 3D Gaussians, each of which has certain properties. Each 3D Gaussian is defined by its mean position $\mu\in\mathbb{R}^{3}$ , covariance matrix $\Sigma$ as the following:
@@
Since the covariance matrix has to be positive semi-definite for differentiable optimization, it is decomposed into a rotation matrix $\mathbf{R}$, expressed in quaternions and a scaling matrix $\mathbf{S}\in\mathbb{R}^{3}$ as:
@@
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM}$ (3)
@@
3.2 Multi-resolution Hash Grid
Encoding neural network inputs into higher dimensional space has been crucial in achieving scene representations of high quality with compact models [50].  [54] furthers fidelity and efficiency by outsourcing the representation power to an auxiliary data structure, namely a multi-resolution hash grid, which allows fast access to learned features at each grid cell on multiple spatial resolutions. Specifically, an input 3D point is first parameterized by a hash code $(\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{l})$, which corresponds to the corner indices of a set of grids with different spatial resolution $\{V_{1},\ldots,V_{L}\}$. Given those indices, feature vectors are obtained and linearly interpolated according to the relative position of $\boldsymbol{x}$ across all $L$ spatial resolutions. Formally, the final multi-resolution hash encoding of a 3D point $\boldsymbol{x}$ becomes $\gamma(\boldsymbol{x})=(\gamma_{1}(\boldsymbol{x}_{1}),\ldots,\gamma_{l}(\boldsymbol{x}_{l}))$, where $\gamma_{l}(\boldsymbol{x}_{l})$ is a feature vector obtained by means of interpolation of hash entries at the grid corners of the $l$th resolution level.
Multi-resolution hash encoding is independent at each level and highly adaptive to the desired finest resolution. Thus, it enables faster convergence during training, does not impede inference speed, and promotes a higher level of detail, which is beneficial for the stylization task using 3DGS representation.
@@ 3.3 Stylized NeRF
StylizedNeRF [28] set up a new benchmark in the field of neural scene stylization. Built upon the framework of NeRF and Nerf-W [50, 45], this work makes use of a mutual learning strategy to combine 2D stylization method AdaIN [27] and a pre-trained NeRF architecture. The high-level idea is to predict a new stylized color conditioned on the style image. This is achieved by replacing the color component with a new style module while keeping the density prediction from the pre-trained NeRF constant to maintain geometric consistency. To facilitate mutual learning between 2D-3D stylization, StylizedNeRF makes the decoder of AdaIN trainable for fine-tuning the final results. Inspired by Nerf-W, they also employ a pre-trained VAE that provides learnable latent codes for the conditioned style inputs. Motivated by the positive results of StylizedNeRF, GSS also makes use of pre-trained Gaussians to better represent the geometric details of the complex scenes.
Figure 3: Here we diagramatically show the overall architecture of our pipeline. (a) contains the overview of our method. (b) shows in detail the top-right part of (a) where we have a deeper look into how the 3D color module combines with 3DGS to produce stylized novel views.
@@ 4 Gaussian Splatting in Style
Our goal is to stylize a complex 3D scene. The first step is to learn the scene information from a given collection of images in a multi-view setting. We achieve this by using a 3DGS backbone. For this objective, we take pretrained models of learnt scenes for the next step. The second stage includes learning the scene information conditioned on various style images. During inference time, we generate novel views conditioned on an input style image. In the following section, we elaborate on the various constituents of our pipeline.
2D Stylization Module The first component of our pipeline is a 2D stylization module. It is based on AdaIN [27]. It consists of three components, a pretrained VGG [72] as the encoder, a convolutional decoder, and adaptive instance normalisation layer. For details, please refer to [28]. As shown in Figure 3, we pass the style image and the rendered view from 3DGS through a pretrained VGG to obtain their features. These features are then passed through the adaptive instance normalization layer and the decoder to obtain the 2D styled image of that particular view. This 2D stylized image acts as a guide for our 3D color module, and gives a rough indication as to how the scene should appear from that specific viewpoint. Therefore, this image, represented by $\mathcal{I}_{style2d}$ forms on what we call it as a Guide Loss $\mathcal{L}_{g}$, which is detailed in Section 4.
@@
3D Color Module This module is the second component in our overall architecture. The input to this module are the mean positions $\mu$ of the 3D Gaussians and a latent code representing the style image obtained through the use of a pretrained encoder[65, 28]. The input positions are fed into a multi-resolution hash grid(MHG) [54] while the latent codes are fed into an FC layer before concatenating them together. This resulting tensor is then passed through a tiny MLP [53, 38] to obtain the new RGB color for each input Gaussian. The use of MHG and tiny MLP was motivated by the fact that each scene has a large number of Gaussians, and could also be in the order of millions for large scenes such as those in the TnT[35] dataset. The predicted RGB from the 3D color module is fed into the 3DGS pipeline, and the Gaussians, with their new colors are rendered to generate a 3D consistent stylised view $\mathcal{I}_{gen}$. To preserve important details in the scene, we compare $\mathcal{I}_{gen}$ with the unstyled view $\mathcal{I}_{i}$ as shown in Equation 5. Since our color module is view dependent, it does not affect the runtime during inference time. Since the geometry is conserved, one forward pass through GSS with the conditioned style input allows us to obtain the stylized RGB for any arbitrary viewpoint.
@@
Loss function As shown in Figure 3, we apply the loss functions on the views generated by the 3D color module. Our loss term is composed as a weighted sum of two terms, the guide loss $\mathcal{L}_{g}$ and the content loss $\mathcal{L}_{c}$, elaborated as follows:
$\mathcal{L}_{g}=||\mathcal{I}_{style2d}-\mathcal{I}_{gen}||$ (4)
$\mathcal{L}_{c}=||\mathcal{I}_{gen}-\mathcal{I}_{i}||$ (5)
@@
where $\mathcal{I}_{style2d},\mathcal{I}_{i},\mathcal{I}_{gen}$ are the 2D stylized image from the 2D stylization module, the unstyled view generated from 3D Gaussian splatting, and the stylized view from our pipeline respectively.
Finally, our total loss $\mathcal{L}_{total}$ can be written as:
$\mathcal{L}_{c}=\lambda_{g}\mathcal{L}_{g}+\lambda_{c}\mathcal{L}_{c}$ (6)
where $\lambda_{g}$ and $\lambda_{c}$ are the weights of the guide loss and content loss respectively.

### v2

@@ Abstract
3D scene stylization extends the work of neural style transfer to 3D. A vital challenge in this problem is to maintain the uniformity of the stylized appearance across multiple views. A vast majority of the previous works achieve this by training a 3D model for every stylized image and a set of multi-view images. In contrast, we propose a novel architecture trained on a collection of style images that, at test time, produces real time high-quality stylized novel views. We choose the underlying 3D scene representation for our model as 3D Gaussian splatting. We take the 3D Gaussians and process them using a multi-resolution hash grid and a tiny MLP to obtain stylized views. The MLP is conditioned on different style codes for generalization to different styles during test time. The explicit nature of 3D Gaussians gives us inherent advantages over NeRF-based methods, including geometric consistency and a fast training and rendering regime. This enables our method to be useful for various practical use cases, such as augmented or virtual reality. We demonstrate that our method achieves state-of-the-art performance with superior visual quality on various indoor and outdoor real-world data.
@@
Figure 1: Given multi-view images of a real-world scene, we perform the task of scene stylization. Unlike previous scene stylization approaches, we do not need to fit a scene to each new style. Employing a neural network, conditioned on a style image, allows us to generalize to a variety of styles. Our method can generate 3D consistent scene stylization at approximately $150$ FPS.
Figure 2: The motivation from our work stems from the requirement of a specialized method that, while stylizing a scene, considers the spatial information into account. We show that to generate stylized novel views of a scene, it is insufficient to stylize the rendered views or train a scene representation model on stylized 2D images. It leads to loss of information, such as deformity in the solid truck’s body shown above.
@@ 1 Introduction
Driven by the curiosity of how a particular image would appear if painted by Van Gogh, Picasso and other artists, style transfer has been at the heart of the image processing community for a long time. It can be seen as a challenge of texture transfer. The core idea of neural style transfer is, given a style image with desired aesthetics and a target image with pursued content, to apply the features of a style image such as colors and texture to the content image while keeping information inside the content image intact. Earlier works [13, 81, 37, 12] devised powerful algorithms for texture synthesis. Further advancements in deep learning gave rise to works [16, 31, 27] that approach the problem by suitably trained neural networks. The ability of neural networks to effectively grasp the features of an image at multiple levels of details made them a preferred choice over non-learning based algorithms[16]. Fuelled by the availability of 3D data, this phenomenon of style transfer was extended to the stylization of scenes[28, 26, 43, 87], and is referred to as scene stylization, which is the primary focus of this work.
@@
The ability to control the visual properties of a scene is central to many practical applications such as avatar modeling or scene editing [60, 51, 64, 91] to name a few. With the wide adaptability and increasing interest in various augmented and virtual reality applications, it is not only necessary that these techniques are accurate, but also real-time capable to ensure the best user experience.
@@
Over the past few years, different methods have been proposed to efficiently represent a complex 3D scene. [59, 82, 77, 71] were some methods leveraging an explicit pointcloud representation. The domain witnessed an incredible pace of growth following the pioneering work of Neural Radiance Fields(NeRFs)[50, 49]. NeRFs gained large popularity, primarily due to its ability to represent a complex scene in the weights of a single multi-layer perceptron(MLP), while capturing fine details during novel view synthesis. Nonetheless, the proposed method required long training time and it was slow during inference. Therefore, many variants of NeRFs emerged such as [66, 86, 1, 54, 57, 3] that improved the seminal work significantly. Recently, 3D Gaussian Splatting(3DGS)[33] further pushed the boundaries of novel view synthesis, especially in terms of consistency and speed. 3DGS is, by design, more explicit in terms of storing scene information in the form of 3D Gaussians, with each Gaussian having certain properties attached to it. Additionally, 3DGS renders views in real-time, thus making it highly suitable for practical applications.
@@
In this paper, we tackle the problem of learning to stylize a set of 3D Gaussians given a set of style images, namely scene stylization. Therefore, we call our approach as Gaussian Splatting in Style(GSS). We condition this representation on a style image to obtain stylized views of the complex 3D scene. Utilizing these Gaussians as the backbone of our approach ensures spatial consistency. Due to the proposed design, our model stylizes a scene given any style image and does not require re-training as some of the previous works on scene stylization such as ARF [87]. Furthermore, our approach does not add any overhead to the existing real-time rendering speed of 3DGS, making it a preferred choice for AR/VR applications. We demonstrate the superiority of our method on well-established consistency metrics, showing the best performance in both short-term and long-term measurements. We further support our claims with the qualitative analysis of stylized rendered views comparing our work to existing scene stylization baselines.
To summarise, our contribution in this paper are as follows:
@@
  • We develop a novel state-of-the-art method, GSS, to perform neural scene stylization in real-time based on 3D Gaussian splatting. We are among the first to perform scene stylization using 3D Gaussians.
@@ 2.2 Novel View Synthesis
2.3 Style Transfer and Scene Stylization
@@
Neural style transfer is the task of transferring the style of a source image onto the target(or content) image[16]. It is a complex task as the transfer needs to preserve the information present in the target image. This problem can also be formulated as that of texture transfer, and was tackled in the works of [13, 12, 81, 37] before the widespread adoption of neural networks. Later, the pioneering work of [16] successfully applied the use of neural networks to the task of style transfer. The work made use of a pretrained VGG[72] to extract the semantic information from an image at different hierarchical levels. [16] optimized a loss function composed of two terms, content loss and style loss, which represented the similarity of the generated image with the content image and the style image respectively. Since [16] worked by running an optimization for each target and style image, it was not suitable for real-time applications. This was further improved by [40, 31, 78, 27] by being real-time and providing the flexibility of general style conditioning. Additionally, the works of [44, 30] added the sophistication of monocular depth estimation for preserving the depth information present in the source image. Furthermore, works of [25, 19, 15, 4, 6, 80] extended the work of 2D style transfer to a video sequence by making use of advancements in methods like optical flow[11]. However, [26] and our experiments show that simply extending 2D style transfer to 3D scenes result in visual artifacts such as blurriness and inconsistency across different views, as also depicted in Figure 6.
@@
Stylizing scenes given a style image has gained prominence in recent years, especially with the arrival of AR/VR headsets in the mainstream. One way to distinguish between different methods is based on their way of scene representation. Works of [26, 52, 2] work with a point cloud based representation. [26] for instance, project the style features from 2D into 3D and transform them into that of the style image before re-projecting them back into 2D for getting stylized views. Similarly, [23] performs scene stylization of indoor scenes in the form of a mesh. Recently, radiance fields [50] gained huge popularity due to their ability to generate novel views with a very high quality. Due to this, a lot of works made use of a NeRF backbone to perform scene stylization. One direction of approaching the problem is by focusing to optimize a radiance for each given style image. Methods such as [32, 87, 41, 8, 56, 84, 89, 61, 90] follow this outline. Methods such as [7, 8] make use of a hypernetwork to impart the style image features onto the rendered views of a complex 3D scene. On the other hand, works such as [28, 43] provide a generalizable scene stylization framework, i.e a stylized scene can be generated at inference time given a conditioned style image input. Our work also follows in this category. Another direction in this domain lies making use of additional information such as depth maps [32] since their primary focus lies in geometry preservation. In this work, instead of a radiance field, we make use of a Gaussian splatting[33] framework.
@@ 3 Preliminaries
3D Gaussian Splatting(3DGS)[33] is an explicit method for scene representation, that is especially useful for high quality real-time rendering. We briefly introduce the main concepts for completeness.
Gaussian Splatting can be considered as a form of point cloud representation. Each scene in 3DGS is represented as a bunch of 3D Gaussians, each having certain properties. Specifically, each 3D Gaussian is defined by its mean position $\mu\in\mathbb{R}^{3}$ and a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as
@@
Since the covariance matrix has to be positive semi-definite for differentiable optimization, it is decomposed into a rotation matrix $\mathbf{R}$ and a scaling matrix $\mathbf{S}\in\mathbb{R}^{3}$ as
@@
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM},$ (3)
@@ 3.2 Stylized NeRF
StylizedNeRF [28] set up a new benchmark in the field of neural scene stylization. Built upon the framework of NeRF and Nerf-W [50, 45], this work makes use of a mutual learning strategy to combine the 2D stylization method AdaIN [27] and a pre-trained NeRF architecture. The high-level idea is to predict a new stylized color conditioned on the style image. This is achieved by replacing the color component with a new style module while keeping the density prediction from the pre-trained NeRF constant to maintain geometric consistency. To facilitate mutual learning between 2D-3D stylization, StylizedNeRF makes the decoder of AdaIN trainable for fine-tuning the final results. Inspired by Nerf-W, they also employ a pre-trained VAE that provides learnable latent codes for the conditioned style inputs. Motivated by the positive results of StylizedNeRF, GSS also makes use of pre-trained Gaussians to better represent the geometric details of the complex scenes.
Figure 3: Here we diagramatically show the overall architecture of our pipeline. We employ a novel 3D Color module that is jointly trained with the 3D Gaussians, to predict the new colors for each Gaussian based on the querying style image at test time.
@@ 4 Gaussian Splatting in Style
Our goal is to stylize a complex 3D scene. Our overall training process is similar to that of 3DGS[33, 38], where the scene is learnt by minimizing the difference between multiple training views and their corresponding ground truth images. In addition to learning the scene representation, we learn a mapping between the different views of the scene and the queried style image. To this end, we employ an MLP that learns the mapping between the various 3D Gaussians and their stylized color based on the style image. During inference time, we generate novel views conditioned on an input style image. It is worth noting that the input style image at inference time may or may not have been a part of the dataset of style images used in the training process. In the following section, we elaborate on the various constituents of our pipeline.
@@
3D Color Module The first component of our pipeline is a 3D Color module, responsible for predicting the colors of each 3D Gaussian based on the conditioning style image. The input to this module are the mean positions $\mu$ of the 3D Gaussians and a latent code representing the style image obtained through the use of a pretrained encoder[65, 28]. The mean input positions are fed into a multi-resolution hash grid(MHG) [54] while the latent codes are fed into an FC layer before concatenating them together. This resulting tensor is then passed through an MLP [53, 38] to obtain a new RGB color for each input Gaussian. The use of MHG and tiny MLP was motivated by the fact that each scene has a large number of Gaussians, and could also be in the order of millions for large scenes such as those in the TnT[35] dataset. The predicted RGB from the 3D color module is fed into the 3DGS pipeline, and the Gaussians, with their new colors are rendered to generate a 3D consistent stylised view $\mathcal{I}_{gen}$. Since our color module is view independent, it does not affect the runtime during inference time. Furthermore, the geometry is conserved, one forward pass through the proposed pipeline with the conditioned style input allows us to obtain the stylized RGB for any arbitrary viewpoint.
2D Stylization Module The 2D Stylization moule is based on AdaIN [27]. It consists of three components, a pretrained VGG [72] as the encoder, a convolutional decoder, and adaptive instance normalisation layer. For details, please refer to [28]. As shown in Figure 3, we pass the style image and the original ground truth view through a pretrained VGG to obtain their features. These features are then passed through the adaptive instance normalization layer and the decoder to obtain the 2D styled image of that particular view. This 2D stylized image acts as a guide for our 3D color module, and gives a rough indication as to how the scene should appear from that specific viewpoint. Therefore, this image, represented by $\mathcal{I}_{style2d}$ forms on what we call it as a Guide Loss $\mathcal{L}_{g}$, which is detailed in Section 4.
Training Regime We build on top of vanilla 3DGS[33]. During the training process, we learn the scene in conjunction with the 3D Color module. To this end, we first start by learning the scene without the color module and train on unstyled training views. The motivation for this ”warm-up” phase arises to encourage the network to learn the original geometry, and therefore have an inherent notion of being consistent.
@@
Loss function As shown in Figure 3, we apply the loss functions on the views generated by the 3D color module. Our loss term is therefore made up of the guide loss $\mathcal{L}_{g}$ and is elaborated as
$\mathcal{L}_{g}=||\mathcal{I}_{style2d}-\mathcal{I}_{gen}||,$ (4)
@@
where $\mathcal{I}_{style2d},\mathcal{I}_{gen}$ are the 2D stylized image from the 2D stylization module, and the stylized view from our pipeline respectively.